# PACP: Priority-Aware Collaborative Perception for Connected and Autonomous Vehicles

### v1

@@ 5.3 Submodular Analysis and Solutions
0:  Output the optimal link establishment matrix $\mathcal{S}$.
@@
In this subsection, we first prove that the objective function of the problem $\mathbf{P_{2}}$ is a submodular function. Specifically, as more CAVs share their perception results, the ego CAV tends to add significant new information for view fusion. However, as the number of CAV increases, each additional CAV provides less new information, i.e., Diminishing Marginal Utility. This concept is crucial in CAVs’ scenarios, ensuring that resources are not wasted by redundant sensors. Moreover, we attempt to adapt a greedy algorithm to solve the above maximization submodular function problem. Besides, we show that a greedy algorithm finds a near-optimal solution with at least $\left(1-e^{-1}\right)$ of the optimal value.
@@
[…]
@@
$\mathbf{P}_{2}:\quad\!$ $\max_{\mathcal{S}}\,\,\sum_{i=1}^{N}{\left(\mathrm{\omega}_{1}\sum_{j=1,j\neq i}^{N}{\mathcal{P}_{ji}s_{ji}d_{ij}^{(n)}}+\mathrm{\omega}_{2}\mathcal{I}\left(\mathcal{S}_{i}\right)\right)}$ (26)
@@ 5.4 Deep Learning-Based Adaptive Compression
(1) Encoding Phase: The input image $i\in\mathbb{R}^{N}$ undergoes transformation to latent representation $h$ (expressed as $h=f(i)$). A quantization function ${Q}$ further refines $h$ into a discrete vector $q\in\mathbb{H}^{D}$, leading to $q=Q(h)$. In V2V scenarios, $q$ is serialized into bitstream $b$, with entropy coding to maximize the bandwidth efficiency.
@@
(2) Decoding Phase: The bitstream $b$ undergoes inverse transformations. This includes dequantization ($\hat{h}=Q^{-1}(q)$) and a reconstruction function ($\hat{i}=H(\hat{h})$) to retrieve the original image. Conventional compression struggles in dynamic V2V networks due to their static nature. The Deep Learning-Based Compression (DBC) addresses this by trainable parameters from training datasets, thus offering adaptability to V2V dynamics. Under DBC architecture, both encoder and decoder utilize convolutional layers. The encoder transforms the input image into a latent representation, $h=f_{\theta}(i)$, with transformation parameters $\theta$ learned from training. The decoder uses a distinct parameter set $\xi$ for reconstruction: $\hat{i}=H_{\xi}(h)$. The training objective is to minimize:
$\mathop{\arg\min}_{\theta,\xi}R\left(b\right)+\beta D(i,\hat{i}),$ (28)
@@
where $R(b)$ and $D(i,\hat{i})$ measure bitrate and distortion using the Mean Square Error (MSE). Constant $\beta$ manages the R-D trade-off. Moreover, we formulate the traditional fixed R-D problem as a multi-R-D problem for adaptability:
$\mathop{\arg\min}_{\theta,\xi}R(\hat{h};\theta)+\beta D\left(i,\hat{i};\theta,\xi\right),$ (29)
@@
Notably, a static $\beta$ can affect reliable decoding and image quality, emphasizing the need to adjust $\beta$ adaptively. In this context, we introduce a DBC mechanism to dynamically modify $\beta$ under dynamic channel conditions (As shown in Fig. 5, given the channel state information (CSI) feedback, we can obtain $\beta$ and compression ratio according to Sec. 5.1). Therefore, the revised problem can be formulated as:
$\mathop{\arg\min}_{\theta,\xi}R(\hat{h};\theta,{r})+G({r})D\left(i,\hat{i};\theta,\xi,{r}\right).$ (30)
@@
It is noted that the adaptive compression network is based on a pre-trained model[25], using historical camera data as training dataset to obtain network parameter $\theta,\xi$ by solving the multi-R-D problem in (30).
@@
(3) Fine-tuning Phase: This adaptive approach enables the ego CAV and nearby CAVs adaptively adjust the data compression ratio according to dynamic channel conditions. As shown in Fig. 5, we introduce a fine-tuning compression strategy to further reduce temporal redundancy between frames. Utilizing a Modulated Autoencoder (MAE), our method first transmits a few uncompressed images to RSU for real-time fine-tuning. As for computational cost, the encoder and decoder of MAE require 0.155 MFLOPs/pixel and 0.241 MFLOPs/pixel [^4] As a metric of computational cost, MFLOPs/pixel denotes the number of million floating point operations performed per pixel for camera perception data., respectively[25]. Considering Tesla FSD as the computational unit, when three CAVs share their perception results at a rate of 10 fps, the average latencies for encoding and decoding processes are observed to be 40.26 ms and 20.89 ms, respectively.
@@
4:     Sort column in descending order and get the indices of the largest $K$ capacity, and set the associated $s_{ij}=1$;
@@
12:  According to Eq.( 30), train encoders (decoders) to compress (reconstruct) raw camera data;
@@ 6.2 Simulation Settings
Our simulations are based on the 3GPP standard[32]. Specifically, the communication range for vehicles is 150 m. Vehicular speeds, unless otherwise stated, ranging from 00 to $50$ km/h (typical speed range on city roads), are generated using the CARLA simulator[27]. Additionally, vehicles are uniformly distributed over a six-lane highway of 200 meters, with three lanes allocated for each traveling direction. To simplify, certain frame types, such as cyclic redundancy checks and Reed-Solomon coding, have been omitted from our simulation. In our simulation setup, we consider a system with $N=10$ cooperative vehicles. Each vehicle has local data of $A_{j}=40$ Mbits and can access up to $K=4$ subchannels. The computation complexity is quantified by $\beta=100$ Cycles/bit, and each vehicle transmits at a power $P_{t}=8$ mW. The CPU capacity, $F_{j}$, of these vehicles varies between 1 GHz and 3 GHz, as highlighted by Lyu et al. in [21]. Additionally, we allocate a bandwidth of $W=200$ MHz and set a power threshold $E_{j}^{T}=1$ kW for each vehicle.
@@ 6.3 Analysis of Experimental Results
In Eq. (13), the utility value $\mathcal{U}_{\mathrm{sum}}$ denotes the weighted utility function of perception quality and perceptual region. As the priority value is determined by the degree of IoU matching, a larger utility value indicates that more data matching the perception results of the ego CAV have been transmitted. Intuitively, the higher the utility value, the higher the accuracy of the perception results of the ego CAV. Fig. 7 presents a comprehensive view of how different communication parameters influence the utility value $\mathcal{U}_{\mathrm{sum}}$. In Fig. 7(a), as the transmission power increases, both PACP and DMDDA show an improvement in the utility value. However, the utility value of PACP consistently surpasses that of DMDDA. Specifically, at a typical power 8.0 mW, PACP is about 8.27% better than DMDDA, underlining the efficiency of the proposed PACP approach over the baseline in terms of better prioritized data communication. Fig. 7(b) indicates that the utility value generally declines as the maximum range increases. Despite this trend, PACP consistently outperforms DMDDA, illustrating its robustness even when the communication range is extended. Fig. 7(c) reveals that as the number of CAVs increases, the utility value increases. After deploying 9 vehicles, the PACP algorithm outstrips DMDDA by a notable margin of 17.79%. This accentuates the scalability and efficacy of PACP in larger vehicular networks. From Fig. 7(d), PACP amplifies the utility value by 8.07% in comparison with DMDDA at a bandwidth of 180 MHz. This suggests that PACP is adept at capitalizing on larger bandwidths to enhance the perception accuracy of the ego CAV.
Intuitively, we can adopt throughput as a metric for cooperative perception. Similar to utility value, a higher throughput often entails a greater amount of valuable information being transmitted. Fig. 8(a) shows that larger power enhances throughput. Fig. 8(b) depicts an inverse correlation between the vehicle distribution maximum range (80 m - 210 m) and throughput. It is noted that the performance of FTS strategy deteriorates because of balancing communication channels of remote vehicles. Remarkably, with five cooperative CAVs, the PACP algorithm augments the average throughput by 20.39% compared with DMDDA. Furthermore, Fig. 8(d) underscores the linear affinity between the network’s bandwidth (80-180 MHz) and throughput. Specifically, a bandwidth of $W=180$ MHz elevates the average throughput by 18.61% using the PACP algorithm against other baselines. To sum up, it is evident that a higher throughput results in a richer amount of high-priority data being transmitted according to Figs. 7-8.
TABLE II: The comparison of PACP with three baselines under different parameters in terms of perception accuracy.
@@
To further analyze the performance of perception accuracy, we conduct extensive experiments to verify the performance of PACP. From Table II, it is evident that the proposed PACP method consistently outperforms other techniques under various settings. PACP, underpinned by a priority-aware perception scheme, demonstrates a remarkable increase in perception accuracy, surpassing FTS by a minimum of 35.38% and DMDDA by 5.71% in scenarios under varying numbers of CAVs. Given a typical bandwidth 200 MHz, the improvement of adapting PACP achieves up to 13.60%. Such performance elevation is attributed to PACP’s adeptness at channel allocation based on CAVs’ significance to the ego vehicle, ensuring optimal transmission of pivotal perception data. Contrarily, fairness-based strategies, as exemplified by FTS’s bandwidth-fairness and DMDDA’s subchannel-fairness, inadvertently lead to suboptimal resource allocation by treating every CAV with uniform significance. This often results in the situation that vital information receives the same resource allocation as non-essential data, consequently diminishing AP@IoU and increasing the chances of missed detections in a BEV context. Not unexpectedly, as the power, the number of CAVs, and bandwidth increase, the amount of information perceived by the ego CAV expands, leading to a consequent rise in the AP@IoU for all perception strategies.

### v2

@@ 5.3 Submodular Analysis and Solutions
10:  Output the optimal link establishment matrix $\mathcal{S}$.
@@
In this subsection, we first prove that the objective function of the problem $\mathbf{P_{2}}$ is a submodular function. The objective function $\mathcal{U}_{\mathrm{sum}}(\mathcal{S},\mathcal{D})$ represents the aggregated utility of the system, incorporating both perception quality and coverage. However, achieving an optimal solution for this objective function is challenging due to its NP-hard nature. As more CAVs share their perception results, the ego CAV gains significant new information for view fusion. However, as the number of CAVs increases, each additional CAV provides less new information, i.e., Diminishing Marginal Utility. This concept is crucial in CAV scenarios to ensure resources are not wasted by redundant sensors. To efficiently solve the submodular function maximization problem, we adapt a greedy algorithm. Submodular functions exhibit the diminishing returns property, allowing a greedy algorithm to find a near-optimal solution. Specifically, we show that a greedy algorithm can achieve at least $\left(1-e^{-1}\right)$ of the optimal value to solve $\mathbf{P_{2}}$ efficiently.
@@
Given that $\mathbf{P}_{2}$ characterizes the link establishment problem and the data rate $\mathcal{D}=\mathcal{D}^{(n)}$ is a constant matrix, the objective function $\mathcal{U}_{\text{sum}}(\mathcal{S},\mathcal{D}^{(n)})$ defined in Eq. (17) exhibits submodularity if $\mathcal{U}_{\text{sum}}(\mathcal{S},\mathcal{D}^{(n)})$ satisfies all the properties outlined in Definition 3 within Sec. 5.2.
@@
$\mathbf{P}_{2}:$ $\max_{\mathcal{S}}\,\,\sum_{i=1}^{N}{\left(\mathrm{\omega}_{1}\sum_{j=1,j\neq i}^{N}{\mathcal{P}_{ji}s_{ji}d_{ij}^{(n)}}+\mathrm{\omega}_{2}\mathcal{I}\left(\mathcal{S}_{i}\right)\right)}$ (30)
@@ 5.4 Deep Learning-Based Adaptive Compression
The Deep Learning-Based Compression (DBC) addresses the challenges by using trainable parameters from training datasets, thus offering adaptability to V2V dynamics. Under DBC architecture, both encoder and decoder utilize convolutional layers. The encoder transforms the input image $\mathbf{x}$ into a latent representation, $\mathbf{z}=f(\mathbf{x};{\theta})$, with transformation parameters $\theta$ learned from training. The decoder uses a distinct parameter set $\xi$ for reconstruction: $\tilde{\mathbf{z}}=H(\mathbf{z};{\xi})$. $\mathbf{\hat{x}}$ is the reconstruction image from the decoder. The training objective is to minimize:
$\mathop{\arg\min}_{\theta,\xi}R\left(\mathbf{\tilde{z}};\theta\right)+\beta D(\mathbf{x},\mathbf{\hat{x}};\theta,\xi),$ (32)
@@
where $R\left(\mathbf{\tilde{z}};\theta\right)=\mathbb{E}\left[-\log_{2}p_{\mathbf{\tilde{z}}}\left(\mathbf{\tilde{z}}\right)\right]$ represents the rate function and $D\left(\mathbf{x},\mathbf{\hat{x}};\theta,\xi\right)=\mathbb{E}\left[\left\|\mathbf{x}-\mathbf{\hat{x}}\right\|^{2}\right]$ denotes the distortion function,where the constant $\beta$ manages the R-D trade-off. Moreover, we formulate the traditional fixed R-D problem as a multi-R-D problem for adaptability:
$\mathop{\arg\min}_{\theta,\xi}R(\hat{h};\theta)+\beta D\left(\mathbf{x},\mathbf{\hat{x}};\theta,\xi\right).$ (33)
@@
It is noted that $\beta$ can affect reliable decoding and image quality, emphasizing the need to adjust $\beta$ adaptively. In this context, we introduce a DBC mechanism to dynamically modify $\beta$ under dynamic channel conditions (As shown in Fig. 7, given the channel state information (CSI) feedback, we can obtain $\beta$ and compression ratio according to Sec. 5.1). Therefore, the revised problem can be formulated as:
$\mathop{\arg\min}_{\theta,\xi}R(\hat{h};\theta,{r})+G({r})D\left(\mathbf{x},\mathbf{\hat{x}};\theta,\xi,{r}\right),$ (34)
@@
where function $G$ is the R-D mapping function, which utilizes a lookup table and converts the compression ratio to the tradeoff parameter $\beta$. Therefore, there is no explicit expression for the function $G$. Moreover, the adaptive compression network is based on a pre-trained model[28], using historical camera data as training dataset to obtain network parameter $\theta,\xi$ by solving the multi-R-D problem in (34).
1) Encoder Phase: The input image $\mathbf{x}$ undergoes transformation to latent representation $\mathbf{z}$. A quantization function ${Q}$ further refines $\mathbf{z}$ into a discrete vector $q\in\mathbb{H}^{D}$, leading to $q=Q(\mathbf{z})$. In V2V scenarios, $q$ is serialized into bitstream $b$, with entropy coding to maximize the bandwidth efficiency. As shown in Fig. 8(a), the encoder compresses the input camera data, reducing its size for transmission while retaining essential information. It comprises the following layers:
  • Convolutional Layers (Conv): There are three convolutional layers with kernel sizes of $9\times 9$, $5\times 5$, and $5\times 5$, and strides of 4, 2, and 2, respectively. These layers progressively reduce the spatial dimensions of the input data from $[H,W,3]$ to $[H/16,W/16,F3]$, where $H$ is the height, $W$ is the width, and $F1$, $F2$, and $F3$ are the number of filters at each layer.
  • Generalized Divisive Normalization (GDN) Layers: GDN layers follow each convolutional layer to normalize the feature maps, promoting sparsity and reducing redundancy, which is critical for efficient compression. The output dimensions remain the same as those of the preceding convolutional layers.
  • Element-wise Modulation: The outputs of the GDN layers are modulated by element-wise multiplication with modulation vectors $\mathbf{m}(h_{ij})=(m_{1}(h_{ij}),m_{2}(h_{ij}),m_{3}(h_{ij}))$ generated by the modulation network. This modulation adapts the compression process to the CSI feedback, $h_{ij}$.
  • Quantization: The modulated data is then quantized, reducing the precision of the latent representation and enabling efficient encoding into a bitstream.
(a) The encoder part.
(b) The decoder part.
Figure 8: The architecture of the deep learning-based adaptive compression.
2) Decoding Phase: The bitstream $b$ undergoes inverse transformations. This includes dequantization ($\hat{\mathbf{z}}=Q^{-1}(q)$) and a reconstruction function ($\hat{\mathbf{x}}=H({\hat{\mathbf{z}}})$) to retrieve the original image. As shown in Fig. 8(b), the decoder reconstructs the original camera data from the compressed bitstream, ensuring that the essential information is preserved and accurately recovered. It consists of:
  • Inverse Generalized Divisive Normalization (IGDN) Layers: IGDN layers, which reverse the normalization effect of the GDN layers, are applied to the quantized latent representation. The output dimensions are $[H/16,W/16,F3]$.
  • Element-wise Demodulation: The outputs of the IGDN layers are demodulated by element-wise multiplication with demodulation vectors $\mathbf{d}(h_{ij})=(d_{1}(h_{ij}),d_{2}(h_{ij}),d_{3}(h_{ij}))$ generated by the demodulation network, adapting the reconstruction to the CSI feedback.
  • Deconvolutional Layers (Deconv): These layers upsample the feature maps to reconstruct the data, with kernel sizes of $5\times 5$, $5\times 5$, and $9\times 9$, and strides of 2, 2, and 4, respectively, restoring the dimensions to $[H,W,3]$.
Figure 9: The architecture of modulation/ demodulation networks.
3) Modulation/Demodulation Networks: As shown in Fig. 9, the modulation and demodulation networks translate the CSI feedback into adaptive modulation and demodulation vectors, respectively, for the encoder and decoder. They include:
  • R-D Mapping: Converts the CSI feedback $h_{ij}$ into rate-distortion lambda parameters.
  • Fully Connected (FC) Layers: Multiple FC layers with ReLU activations introduce non-linearity and enhance the representation capacity of the network.
  • Exponential Layer (E-Layer): Ensures that the output vectors are positive, necessary for the subsequent modulation and demodulation processes.
@@
4) Fine-tuning Phase: This adaptive approach enables the ego CAV and nearby CAVs to adaptively adjust the data compression ratio according to dynamic channel conditions. As shown in Fig. 7, we introduce a fine-tuning compression strategy to further reduce temporal redundancy between frames. Utilizing a Modulated Autoencoder (MAE), our method first transmits a few uncompressed images to RSU for real-time fine-tuning.
Even though digitization and quantization modules are present in the encoder and decoder, we employ scalar quantization by rounding to the nearest neighbor during inference. During training, this is replaced by additive uniform noise as a proxy, i.e., $\mathbf{\tilde{z}}=\mathbf{z}+\Delta\mathbf{z}$ with $\Delta\mathbf{z}\sim\mathcal{U}(-\frac{1}{2},\frac{1}{2})$. This approach allows the network to approximate gradients during the backward pass, enabling effective backpropagation through these non-differentiable layers. This technique ensures that the network can be trained end-to-end, despite the presence of quantization layers.
As for computational cost, the encoder and decoder of MAE require 0.155 MFLOPs/pixel and 0.241 MFLOPs/pixel [^5] As a metric of computational cost, MFLOPs/pixel denotes the number of million floating point operations performed per pixel for camera perception data., respectively[28]. Considering Tesla FSD as the computational unit, when three CAVs share their perception results at a rate of 10 fps, the average latencies for encoding and decoding processes are observed to be 40.26 ms and 20.89 ms, respectively.
@@
4:     Sort column of link establishment decision matrix in descending order and get the indices of the largest $K$ capacity, and set the associated $s_{ij}=1$;
@@
12:  Train encoders and decoders to compress and reconstruct raw camera data according to Eq. (34);
@@ 6.2 Simulation Settings
Our simulations are based on the 3GPP standard[39]. Specifically, the communication range for vehicles is 150 m. Vehicular speeds, unless otherwise stated, ranging from 00 to $50$ km/h (typical speed range on city roads), are generated using the CARLA simulator[34]. Additionally, vehicles are uniformly distributed over a six-lane highway of 200 meters, with three lanes allocated for each traveling direction. To simplify, certain frame types, such as cyclic redundancy checks and Reed-Solomon coding, have been omitted from our simulation. Table III presents the main simulation parameters utilized in our research.
TABLE III: Simulation Parameters
Parameter Value
Number of vehicles ($N$) 10
Local data per vehicle ($A_{j}$) 40 Mbits
Number of subchannels ($K$) 4
Computation complexity ($\beta$) 100 Cycles/bit
Transmission power ($P_{t}$) 8 mW
CPU capacity ($F_{j}$) 1 GHz - 3 GHz
Bandwidth ($W$) 200 MHz
Power threshold ($E_{j}^{T}$) 1 kW
Weight for perception quality ($w_{1}$) $1\times 10^{-2}$
Weight for perceptual region ($w_{2}$) $1\times 10^{-3}$
Parameter ($\eta$) 1
Compression ratio range ($r_{\min},r_{\max}$) (0.3, 0.95)
@@ 6.3 Analysis of Experimental Results
In Eq. (17), the utility value $\mathcal{U}_{\mathrm{sum}}$ denotes the weighted utility function of perception quality and perceptual region. As the priority value is determined by the degree of IoU matching, a larger utility value indicates that more data matching the perception results of the ego CAV have been transmitted. Intuitively, the higher the utility value, the higher the accuracy of the perception results of the ego CAV. Fig. 11 presents a comprehensive view of how different communication parameters influence the utility value $\mathcal{U}_{\mathrm{sum}}$. In Fig. 11(a), as the transmission power increases, both PACP and DMDDA show an improvement in the utility value. However, the utility value of PACP consistently surpasses that of DMDDA. Specifically, at a typical power 8.0 mW, PACP is about 8.27% better than DMDDA, underlining the efficiency of the proposed PACP approach over the baseline in terms of better prioritized data communication. Fig. 11(b) indicates that the utility value generally declines as the maximum range increases. Despite this trend, PACP consistently outperforms DMDDA, illustrating its robustness even when the communication range is extended. Fig. 11(c) reveals that as the number of CAVs increases, the utility value increases. After deploying 9 vehicles, the PACP algorithm outstrips DMDDA by a notable margin of 17.79%. This accentuates the scalability and efficacy of PACP in larger vehicular networks. From Fig. 11(d), PACP amplifies the utility value by 8.07% in comparison with DMDDA at a bandwidth of 180 MHz. This suggests that PACP is adept at capitalizing on larger bandwidths to enhance the perception accuracy of the ego CAV. In Fig. 14, as the energy consumption constraints $E^{T}$ increase, both PACP and DMDDA show an improvement in the utility value. However, the utility value of PACP consistently surpasses that of DMDDA. Specifically, at a typical constraint of 10 W, PACP achieves about 21.34% higher utility value than DMDDA.
Intuitively, we can adopt throughput as a metric for cooperative perception. Similar to utility value, a higher throughput often entails a greater amount of valuable information being transmitted. Fig. 12(a) shows that larger power enhances throughput. Fig. 12(b) depicts an inverse correlation between the vehicle distribution maximum range (80 m - 210 m) and throughput. It is noted that the performance of FTS strategy deteriorates because of balancing communication channels of remote vehicles. Remarkably, with five cooperative CAVs, the PACP algorithm augments the average throughput by 20.39% compared with DMDDA. Furthermore, Fig. 12(d) underscores the linear affinity between the network’s bandwidth (80-180 MHz) and throughput. Specifically, a bandwidth of $W=180$ MHz elevates the average throughput by 18.61% using the PACP algorithm against other baselines. To sum up, it is evident that a higher throughput results in a richer amount of high-priority data being transmitted according to Figs. 11-12.
TABLE IV: The comparison of PACP with three baselines under different parameters in terms of perception accuracy.
@@
Figure 14: The Utility value vs Energy consumption constraints $E^{T}$ (W).
To further analyze the performance of perception accuracy, we conduct extensive experiments to verify the performance of PACP. From Table IV, it is evident that the proposed PACP method consistently outperforms other techniques under various settings. PACP, underpinned by a priority-aware perception scheme, demonstrates a remarkable increase in perception accuracy, surpassing FTS by a minimum of 35.38% and DMDDA by 5.71% in scenarios under varying numbers of CAVs. Given a typical bandwidth 200 MHz, the improvement of adapting PACP achieves up to 13.60%. Such performance elevation is attributed to PACP’s adeptness at channel allocation based on CAVs’ significance to the ego vehicle, ensuring optimal transmission of pivotal perception data. Contrarily, fairness-based strategies, as exemplified by FTS’s bandwidth-fairness and DMDDA’s subchannel-fairness, inadvertently lead to suboptimal resource allocation by treating every CAV with uniform significance. This often results in the situation that vital information receives the same resource allocation as non-essential data, consequently diminishing AP@IoU and increasing the chances of missed detections in a BEV context. Not unexpectedly, as the power, the number of CAVs, and bandwidth increase, the amount of information perceived by the ego CAV expands, leading to a consequent rise in the AP@IoU for all perception strategies.